OpenStudy (anonymous):

For the function f(x) = –2(x + 3)2 − 1, identify the vertex, domain, and range.

OpenStudy (amoodarya):

\[f(x) = –2(x + 3)^2 − 1\] domain ; no denominator ,no radical ---->\[D_f=R\] range: \[f(x) = –2(x + 3)^2 − 1\\y=–2(x + 3)^2 − 1\\y+1=-2(x+3)^2\\(x+3)^2=\frac{y+1}{-2}\\(x+3)^2 \geq 0 \rightarrow \frac{y+1}{-2} \geq 0\\ \rightarrow y+1 \leq 0 \rightarrow y \leq -1\] for vertex \[y=a(x+b)^2 +c \rightarrow x_{vertex}=-b\]

OpenStudy (anonymous):

so would the vertex be (-3, -1) or (3, -1) ?

OpenStudy (anonymous):

-3 right?

OpenStudy (anonymous):

okay yeah i got it right thank you!